OpenStudy (anonymous):

You earn $245 in gross pay a week and 2/5 of your check is taken out for taxes and deductions. How much of your check is taken out for taxes and deductions?

OpenStudy (anonymous):

A. $49 B. $98 C. $150

OpenStudy (anonymous):

\[$ 245 \times \frac{2}{5} = $ 49 \times 2 = $ 98\]

OpenStudy (anonymous):

thx.... close call 0.0